OpenStudy (anonymous):

Two concentric spheres have radii of 5" and 6". Find the volume of the space between them.

OpenStudy (anonymous):

\[4/3 \pi (6^{3}-5^{3})\] \[4/3 \pi (216-125)\] \[4/3 \pi(91)\] \[364/3 \pi\]